Please include the following in your work on each word problem:
a) A statement defining any unknown quantities in the problem in terms of variables.
b) An equation involving the variable(s) corresponding to the information given in the problem.
c) A complete solution to the equation, showing all work.
d) A complete sentence that answers the question asked, including appropriate units.


The number of aluminum cans used each year varies directly as the number of people using the cans. If 225 people use 60,000 cans in one year, how many cans are used each year in Nashville, which has a population of 691,243?


Answer by rothauserc(4718)

let c be the number of cans people used, p be the number of people and k be the constant of proportionality
:
a) the number of cans used by people equals the product of the number of people and the constant of proportionality
:
b) c = p * k
:
c) 60000 = 225 * k
:
k = 60000/225 = 4000/15 = 800/3
:
c = 691243 * (800/3) = 184331466.67 is approximately 184,331,467
:
d) The population of people in Nashville used 184,331,467 cans in a year
: